In rising from rest to rest in 8 seconds, a lift accelerates uniformly to its maximum speed and then retards uniformly. The retardation is a third of the acceleration and the distance travelled is 20m. Find the acceleration, retardation and the maximum speed reached


1
Expert's answer
2021-02-24T12:49:18-0500

Let the maximum speed attained by the lift = "v_{1}" m/s

Let the upward acceleration of the lift = "a" m/s2

Net upward acceleration = "(a-g)" m/s2

Retardation = "\dfrac{1}{3}a" m/s2

Net retardation = "-(\dfrac{1}{3}a+g)" m/s2

Total time taken by the lift = "8" s

Total distance travelled = "20" m

During upward acceleration :

"v=u+at_{1}"

"\Rightarrow v_{1}=0+(a-g)t_{1}"

"\Rightarrow t_{1}=\dfrac{v_{1}}{a-g}" (1)


"v^2-u^2=2as_{1}"

"\Rightarrow v_{1}^2-0=2(a-g)s_1"

"\Rightarrow s_1=\dfrac{v_1^2}{2(a-g)}" (2)


During retardation :

"v=u+at_{2}"

"\Rightarrow 0=v_{1}-(\dfrac{1}{3}a+g)t_{2}"

"\Rightarrow t_{2}=\dfrac{3v_{1}}{a-3g}" (3)


"v^2-u^2=2as_2"

"\Rightarrow 0-v_1^2=-2(\dfrac{1}{3}a+g)s_2"

"\Rightarrow s_2=\dfrac{3v_1^2}{2(a+3g)}" (4)


Total time taken,

"t_{1}+t_{2}=8 s"

"\Rightarrow \dfrac{v_{1}}{a-g}+\dfrac{2v_{1}}{a-3g}=8"

"\Rightarrow v_{1}a=2a^2+4ag-6g^2" (5)


Total Distance covered,

"s_1+s_2=20m"

"\Rightarrow \dfrac{v_1^2}{2(a-g)}+\dfrac{3v_1^2}{2(a+3g)}=20"

"\Rightarrow v_1^2a=10a^2+20ag-20g^2" (6)


Multiplying eq.(5) by 5 and subtracting from eq.(6), we get

"\Rightarrow v_1^2a-5v_1a=0"

"\Rightarrow v_1a(v_1-5)=0"

Since "v_1a\cancel{=}0",

"\Rightarrow v_1-5=0"

"\Rightarrow v_1=5ms^{-1}"


Similarly we can find out the value of acceleration by putting the value of "v_1" in eq.(5)

"\Rightarrow a=10.45ms^{-2}"

Retardation "=-\dfrac{1}{3}a=-3.48ms^{-2}"